# Accreditation Against Limited Adversarial Noise

### v3

@@ III.1 Trap and Target Circuits
The above assumptions have mentioned twirling CPTP error to stochastic Pauli error, which I define formally in Def. 13. This is an important step in all accreditation protocols, as it reduces the error to a known, more easily quantified form.
@@
$\bullet$ Two algorithms, $P_{\textit{trap}}$ and $P_{\textit{targ}}$, that generate the required trap and target circuits, respectively.
@@
I note that the required (by Algorithm 2) single-qubit unitaries, $\mathcal{U}_{M}$, will exist for any single-qubit measurement and that the outcome of the circuit – if the single-qubit gates of the circuit are not known, as is the case for Bob as they are redacted for him – is irretrievable from the measurement outcomes without the key. This hiding of the circuit outputs is information-theoretically secure (i.e. has perfect security) [50], assuming the single-qubit gates are redacted, and comparable to the one-time pads in universal blind quantum computing [51]. Henceforth, I will denote $P_{\textit{targ}}$ and $P_{\textit{targ}}$, with the changes implemented by Algorithm 2 to hide their outputs, by $P_{\textit{targ}}^{\prime}$ and $P_{\textit{targ}}^{\prime}$ respectively. I note that if you have the key, the output of the circuit can be easily recovered by XOR-ing each measurement outcome with the corresponding bit in the key.
@@ III.2 Presentation of the Upgraded Accreditation Protocol
However, before the full protocol can be presented, in Protocol 3, I must establish the statistical foundations of the new protocol – in Sec. III.2.1 – and the core mechanics of the protocol – in Sec. III.2.2. These statistical methods will be used to evaluate the $P_{0}\in[0,1]$ that defines Bob’s particular choice of SPSCLβ (as in Def. 12), using multiple trap executions ( all within a single use of Protocol 1). If a target circuit is then executed within that same single use of Protocol 1 as the trap circuit executions, this allows the probability of error (as in Def. 11) of the execution of the target circuit to be bounded upper (using that all circuit executions in a single use of Protocol 1 have error within a single SPSCLβ). Due to the argument in Ref. [23, Appendix Sec. 1], this, in turn, upper bounds the ideal-actual variation distance (as defined in Def. 1) of the target circuit execution.
@@ III.2.1 Statistical Basis of the New Accreditation Protocol
Before presenting the accreditation protocol, it is first useful to present a purely statistical lemma (Lemma 3) that will later enable the accreditation protocol. This presentation first begins with Def. 14.
Definition 14.
For any set of real values, =$\mathcal{R}=\big{\{}r_{j}\in\mathbb{R}^{+}\big{\}}^{|\mathcal{R}|}_{j=1}$, define:
$\underline{\mathrm{Avg}\big{(}\mathcal{R}\big{)}}=\dfrac{1}{|\mathcal{R}|}\sum_{j=1}^{|\mathcal{R}|}\bigg{(}r_{j}\bigg{)}.$ (7)
Later the set $\mathcal{R}$ will denote set of the respective probabilities of error in each trap circuit execution within a single use of Protocol 1 but for the duration of Lemma 3 $\mathcal{R}$ is just considered to contain positive real numbers with no meaning attached to them.
@@
$y\leq\bigg{(}1+2\beta\bigg{)}\mathrm{Avg}(\mathcal{R}).$ (8)
@@
Let =$\mathcal{R}=\{r_{j}\in\mathbb{R}^{+}\text{ }|\text{ }1\leq j\leq|\mathcal{R}|\}$. By assumption,
$\forall r_{j}$ $\in\mathcal{R},r_{j}\in[(1-\beta)P_{0},(1+\beta)P_{0}]$ (9)
$\Rightarrow\forall r_{j}$ $\in\mathcal{R},(1-\beta)P_{0}\leq r_{j}.$ (10)
Therefore,
$(1-\beta)P_{0}$ $=\dfrac{1}{|\mathcal{R}|}\sum_{j=1}^{|\mathcal{R}|}\bigg{(}(1-\beta)P_{0}\bigg{)}\leq\dfrac{1}{|\mathcal{R}|}\sum_{j=1}^{|\mathcal{R}|}\bigg{(}r_{j}\bigg{)}=\mathrm{Avg}(\mathcal{R})$ (11)
$\Rightarrow P_{0}$ $\leq\dfrac{\mathrm{Avg}(\mathcal{R})}{1-\beta}.$ (12)
Then, as $\forall y\in[(1-\beta)P_{0},(1+\beta)P_{0}]$, $y\leq(1+\beta)P_{0}$; using Eqn. 12, I conclude that $\forall y\in[(1-\beta)P_{0},(1+\beta)P_{0}]$,
$y\leq(1+\beta)P_{0}\leq\dfrac{1+\beta}{1-\beta}\mathrm{Avg}(\mathcal{R}).$ (13)
Using the Taylor series of $1/(1-\beta)$ (which converges for all values $\beta$ can take) and neglecting quadratic (or higher) terms:
$\dfrac{1+\beta}{1-\beta}\approx 1+2\beta\Rightarrow y\leq\bigg{(}1+2\beta\bigg{)}\mathrm{Avg}(\mathcal{R}),$ (14)
where I have assumed that the approximation from neglecting higher order terms in the Taylor series is tight enough to not affect the inequality. ∎
@@ III.2.2 Core Mechanics of the New Accreditation Protocol
the probability of error of any execution of a circuit where all error is Pauli twirl, during the same single use of Protocol 1 as the trap circuit executions can be upper bounded by:
$\dfrac{1+2\beta}{k}\big{(}\bar{v}+\theta\big{)},$ (15)
@@
$\bullet$ $\alpha\in[0,1]$ is the confidence required of the bound in Eqn. 15.
@@
$p_{j}\leq\bigg{(}1+2\beta\bigg{)}\mathrm{Avg}(\mathcal{R}).$ (17)
The probability a specific trap returns an flags/detects that an error occurs by returning an incorrect measurement outcome, given error occurs, is (as assumed in the lemma statement) lower bounded by $k\in[0,1]$.
Therefore, by also using the independence between circuit executions of whether error occurs, $\mathrm{Avg}(\mathcal{R})$ can be approximated, by checking if the trap circuit executions detect error, – to within additive error, $\theta$, with confidence $\alpha$ – using Hoeffding’s inequality [52]. This only requires that $|\mathcal{R}|\geq N_{\mathrm{Tr}}=\bigg{\lceil}\dfrac{2}{\theta^{2}}\ln{\bigg{(}\dfrac{2}{1-\alpha}\bigg{)}}\bigg{\rceil}+1$, to provide enough trap circuit executions that the approximation of $\mathrm{Avg}(\mathcal{R})$ is within the required error, $\theta$.
Let $\bar{\nu}$ denote the experimentally obtained approximation of the probability a randomly selected (from $\mathcal{R}$) trap returns an output implying error has occurred (i.e. the fraction of trap circuit executions returning an “incorrect” measurement outcome in a single execution of Protocol 1).
@@
$p_{j}\leq\dfrac{1+2\beta}{k}\big{(}\bar{v}+\theta\big{)}.$ (18)
@@ III.2.3 Formal Presentation of the Accreditation Protocol and Proof of its Correctness
$\mathbb{Return}:$ $\mathrm{TargResult}$, $\frac{1+2\beta}{k}\big{(}\mathrm{TrapResult}+\theta\big{)}$.
@@
The correct functioning of Algorithm 3, as an accreditation protocol, in the adapted/modified problems setting is proven in Theorem 1.
@@
In the adapted cryptographic setting (as defined in Sec. II); given efficient classical algorithms, $P^{\prime}_{\textit{targ}}$ and $P^{\prime}_{\textit{trap}}$, that generate trap and target circuits, respectively (as in Sec. III.1); Protocol 3 allows Alice to execute a circuit equivalent (in output distribution, when no error occurs) to any sampBQP circuit, $\mathcal{C}$, and obtain a bound on the ideal-actual variation distance of that execution (to within arbitrary accuracy $\theta\in\mathbb{R}^{+}$, with arbitrary confidence, $\alpha\in[0,1]$).
@@
$p_{t}\leq\bigg{(}\dfrac{1+2\beta}{k}\bigg{)}\big{(}\bar{v}+\theta\big{)},$ (19)
@@
As mentioned, for Eqn. 19 to hold, it is required that the probability of error in each trap is independent of the outcomes of every prior trap. This is not initially guaranteed as Bob is free to choose probabilities based on previous measurement outcomes.
@@
Additionally, due to the $N_{tt}\in\mathbb{N}_{0}$ extra trap circuits executed during Protocol 1, Bob does not know what he has done to the previous circuits as he has no idea which are “decoy” circuit executions and will be discarded (so he cannot know what error he has applied to the “true” circuit executions). In fact, as Bob is not told $\theta,\alpha$, or $N_{tt}$, he cannot even calculate the various probabilities that he has applied each possible combination of errors – or lack thereof – to to the “true” trap circuit executions.
@@
With the probability of error in any target circuit, $\mathcal{C}_{\text{targ}}$, executed within the same single use of Protocol 1 bounded – as in Eqn. 19 – a bound on the ideal-actual variation distance of the execution of that target circuit, $\nu\big{[}\tilde{\mathcal{C}}_{\text{targ}}\big{]}$, can be obtained. By following Ref. [23, Appendix Sec. 1], I derive that the ideal-actual variation distance of the execution of that target circuit is upper bounded by the probability of error occurring in that execution. Therefore,
$\nu\big{[}\tilde{\mathcal{C}}_{\text{targ}}\big{]}\leq\dfrac{1+2\beta}{k}\big{(}\bar{v}+\theta\big{)}.$ (20)
@@ IV Discussion
In this paper I have upgraded the accreditation protocol of Ref. [23] to consider all noise/error to be adversarial. This has necessitated the development of a new model of adversarial error (i.e. the adapted cryptographic setting presented in Sec. II), where Bob is limited according to experimentally derived rules. These rules (that single-qubit gates are hidden from Bob and that the probability of error in similar circuits executed in quick succession have similar probabilities of error), encapsulated in Protocol 1, allow my upgraded accreditation protocol to function almost identically when experiencing adversarial error as the one in Ref. [22] does when the error is IID. Therefore, the desirable qualities the protocol from Ref. [22], such as not requiring lengthening the circuit to accredit, or adding excessively many extra single-qubit gates, or any extra two-qubit gates, or ancilla qubits, or trusting any aspect of a computation, or extra connectivity between qubits are preserved; leading to no diminution in its suitability for near-term usage.
@@
A slight improvement can be obtained using Ref. [23, Theorem 1], which allows the error in single-qubit gates to be weakly gate dependent: this could be added to my cryptographic model by allowing Bob to define error in the – still redacted – single-qubit gates that depends on the single-qubit gates [^11] Single-qubit gates would still have to be redacted to prevent Bob identifying the target but he could, perhaps, specify some dependence without seeing the single-qubit gates e.g. “if that gate is an $\hat{X}$ gate…”. but the differing CPTP maps – for each single-qubit gate – must differ, in the diamond norm, by at most some small known value. This is a further step towards reality and the restriction on the differences between the gate-dependent error reflects the very small errors in single-qubit gates [34, Fig.5].
V Acknowledgements
I would like to thank Animesh Datta and Theodoros Kapourniotis for useful conversations. This work was supported, in part, by an EPSRC IAA grant (G.PXAD.0702.EXP), the UKRI ExCALIBUR project QEVEC (EP/W00772X/2), and the Quantum Advantage Pathfinder (EP/X026167/1).

### v4

@@ III.1 Trap and Target Circuits
The above assumptions have mentioned twirling CPTP error to stochastic Pauli error, which I define formally in Def. 9. This is an important step in all accreditation protocols, as it reduces the error to a known, more easily quantified form.
@@
$\bullet$ Two algorithms, $P_{\textit{trap}}$ and $P_{\textit{targ}}$, that generate trap and target circuits, respectively, without hidden outputs.
@@
I note that the required (by Algorithm 2) single-qubit unitaries, $\mathcal{U}_{M}$, will exist for any single-qubit measurement and that the outcome of the circuit – if the single-qubit gates of the circuit are not known, as is the case for Bob as they are redacted for him – is irretrievable from the measurement outcomes without the key. This hiding of the circuit outputs is information-theoretically secure (i.e. has perfect security) [53], assuming the single-qubit gates are redacted, and comparable to the one-time pads in universal blind quantum computing [54]. Henceforth, I will denote $P_{\textit{targ}}$ and $P_{\textit{targ}}$, with the changes implemented by Algorithm 2 to hide their outputs, by $P_{\textit{targ}}^{\prime}$ and $P_{\textit{targ}}^{\prime}$ respectively. I note that if you have the key, the output of the circuit can be easily recovered by XOR-ing each measurement outcome with the corresponding bit in the key. I will additionally use “outputs” to refer to the results of traps and targets after the key has been used to undo the effect of hiding the outputs from Bob.
@@ III.2 Presentation of the Upgraded Accreditation Protocol
However, before the full protocol can be presented, in Protocol 3, I must establish the statistical foundations of the new protocol – in Sec. III.2.1 – and the core mechanics of the protocol – in Sec. III.2.2. These statistical methods will be used to evaluate the $P_{0}\in[0,1]$ that defines Bob’s particular choice of SPSCLβ (as in Def. 8), using multiple trap executions ( all within a single use of Protocol 1). If a target circuit is then executed within that same single use of Protocol 1 as the trap circuit executions, this allows the probability of error (as in Def. 7) of the execution of the target circuit to be bounded upper (using that all circuit executions in a single use of Protocol 1 have error within a single SPSCLβ). Due to the argument in Ref. [26, Appendix Sec. 1], this, in turn, upper bounds the ideal-actual variation distance (as defined in Def. 1) of the target circuit execution.
@@ III.2.1 Statistical Basis of the New Accreditation Protocol
Before presenting the accreditation protocol, it is first useful to present a purely statistical lemma (Lemma 3) that will later enable the accreditation protocol.
For any set of real values, $\mathcal{R}=\big\{r_{j}\in\mathbb{R}^{+}\big\}^{|\mathcal{R}|}_{j=1}$, let $\mathrm{Avg}\big(\mathcal{R}\big)$ denote the average of $\mathcal{R}$. Later the set $\mathcal{R}$ will denote set of the respective probabilities of error in each trap circuit execution within a single use of Protocol 1 but for Lemma 3 $\mathcal{R}$ is just considered to contain positive real numbers with no meaning attached to them.
@@
$y\leq\frac{1+\beta}{1-\beta}\mathrm{Avg}(\mathcal{R}).$ (5)
@@
Assume $y$ takes the max possible value, $(1+\beta)P_{0}$, and $\mathrm{Avg}(\mathcal{R})$ takes the min possible value, $(1-\beta)P_{0}$. The claimed inequality can then be checked to hold. ∎
@@ III.2.2 Core Mechanics of the New Accreditation Protocol
the probability of error of any execution of a circuit where all error is Pauli twirled, during the same single use of Protocol 1 as the trap circuit executions can be upper bounded by:
$\dfrac{1+\beta}{k(1-\beta)}\big(\bar{v}+\theta\big),$ (6)
@@
$\bullet$ $\alpha\in[0,1]$ is the confidence required of the bound in Eqn. 6.
@@
$p_{j}\leq\frac{1+\beta}{1-\beta}\mathrm{Avg}(\mathcal{R}).$ (8)
The probability a specific trap returns detects that an error occurs by returning an incorrect measurement outcome, given error occurs, is (as assumed in the lemma statement) lower bounded by $k\in[0,1]$.
Therefore, by also using the independence between circuit executions of whether error occurs, $\mathrm{Avg}(\mathcal{R})$ can be approximated, by checking if the trap circuit executions detect error, – to within additive error, $\theta$, with confidence $\alpha$ – using Hoeffding’s inequality [55]. This only requires that $|\mathcal{R}|\geq N_{\mathrm{Tr}}=\bigg\lceil\dfrac{2}{\theta^{2}}\ln{\bigg(\dfrac{2}{1-\alpha}\bigg)}\bigg\rceil+1$, to provide enough trap circuit executions that the approximation – acquired by sampling – of $\mathrm{Avg}(\mathcal{R})$ is within the required error, $\theta$.
Let $\bar{\nu}$ denote the experimentally obtained approximation of the probability a uniformly selected (from $\mathcal{R}$) trap returns an output implying error has occurred (i.e. the fraction of trap circuit executions returning an “incorrect” measurement outcome in a single execution of Protocol 1).
@@
$p_{j}\leq\dfrac{1+\beta}{k(1-\beta)}\big(\bar{v}+\theta\big).$ (9)
@@ III.2.3 Formal Presentation of the Accreditation Protocol and Proof of its Correctness
$\mathbb{Return}:$ $\mathrm{TargResult}$, $\frac{1+\beta}{k(1-\beta)}\big(\mathrm{TrapResult}+\theta\big)$.
@@
The correct functioning of Algorithm 3, as an accreditation protocol, in the adapted/modified problem setting is proven in Theorem 1.
@@
In the adapted cryptographic setting (as defined in Sec. II); given efficient classical algorithms, $P^{\prime}_{\textit{targ}}$ and $P^{\prime}_{\textit{trap}}$, that generate trap and target circuits, respectively (as in Sec. III.1); Protocol 3 allows Alice to execute a circuit that is equivalent (in output distribution, when no error occurs) to any given sampBQP circuit, $\mathcal{C}$, and obtain a bound on the ideal-actual variation distance of that execution (to within arbitrary accuracy $\theta\in\mathbb{R}^{+}$, with arbitrary confidence, $\alpha\in[0,1]$).
@@
$p_{t}\leq\bigg(\dfrac{1+\beta}{k(1-\beta)}\bigg)\big(\bar{v}+\theta\big),$ (10)
@@
As mentioned, for Eqn. 10 to hold, it is required that the probability of error in each trap is independent of the outcomes of every prior trap. This is not initially guaranteed as Bob is free to choose probabilities based on previous measurement outcomes.
@@
Additionally, due to the $N_{tt}\in\mathbb{N}_{0}$ extra trap circuits executed during Protocol 1, Bob does not know what he has done to the previous circuits as he has no idea which are “decoy” circuit executions and will be discarded (so he cannot know what error he has applied to the “true” circuit executions). In fact, as Bob is not told $\theta,\alpha$, or $N_{tt}$, he cannot even calculate the various probabilities that he has applied each possible combination of errors – or lack thereof – to the “true” trap circuit executions.
@@
With the probability of error in any target circuit, $\mathcal{C}_{\text{targ}}$, executed within the same single use of Protocol 1 bounded – as in Eqn. 10 – a bound on the ideal-actual variation distance of the execution of that target circuit, $\nu\big[\tilde{\mathcal{C}}_{\text{targ}}\big]$, can be obtained. By following Ref. [26, Appendix Sec. 1], I derive that the ideal-actual variation distance of the execution of that target circuit is upper bounded by the probability of error occurring in that execution. Therefore,
$\nu\big[\tilde{\mathcal{C}}_{\text{targ}}\big]\leq\dfrac{1+\beta}{k(1-\beta)}\big(\bar{v}+\theta\big).$ (11)
@@ IV Discussion
In this paper, I have upgraded the accreditation protocol of Ref. [26] to consider all noise/error to be adversarial. This has necessitated the development of a new model of adversarial error (i.e. the adapted cryptographic setting presented in Sec. II), where Bob is limited according to experimentally derived rules. These rules (that single-qubit gates are hidden from Bob and that the probability of error in similar circuits executed in quick succession have similar probabilities of error), encapsulated in Protocol 1, allow my upgraded accreditation protocol to function almost identically when experiencing adversarial error as the one in Ref. [25] does when the error is IID. Therefore, the desirable qualities the protocol from Ref. [25], such as not requiring lengthening the circuit to accredit, or adding excessively many extra single-qubit gates, or any extra two-qubit gates, or ancilla qubits, or trusting any aspect of a computation, or extra connectivity between qubits are preserved; leading to no diminution in its suitability for near-term usage.
@@
A slight improvement can be obtained using Ref. [26, Theorem 1], which allows the error in single-qubit gates to be weakly gate dependent: this could be added to my cryptographic model by allowing Bob to define error in the – still redacted – single-qubit gates that depends on the single-qubit gates [^10] Single-qubit gates would still have to be redacted to prevent Bob identifying the target but he could, perhaps, specify some dependence without seeing the single-qubit gates e.g. “if that gate is an $\hat{X}$ gate…”. but the differing CPTP maps – for each single-qubit gate – must differ, in the diamond norm, by at most some small known value. This is a further step towards reality and the restriction on the differences between the gate-dependent error reflects the very small errors in single-qubit gates [37, Fig.5].
V Acknowledgements and Declarations
I would like to thank Animesh Datta and Theodoros Kapourniotis for useful conversations. I would also like to thank an anonymous reviewer for suggesting a shorter proof of Lemma 3. This work was supported, in part, by an EPSRC IAA grant (G.PXAD.0702.EXP), the UKRI ExCALIBUR project QEVEC (EP/W00772X/2), and the Quantum Advantage Pathfinder (EP/X026167/1).
The author declares no conflicts of interest